Assume a corporation has earnings before depreciation and taxes of $100,000, depreciation of $40,000, and a 24 percent tax bracket a. Compute its cash flow using the following format. Earnings before depreciation and taxes Deprociation Eamings beflore taxes Taxes Earnings aftor taxes Depreciation Cash Soww b. Compute the cash flow for the company if depreciation is only $20,000. All other factors are the same. Cash flow
Earrings before depreciation and taxes Earnings before taxes Taxes Earmings after taxes 166 points Cash Bow Hnt b. Compute the cash flow for the c ompany if depreclation is only $20,000. All other factors are the same. Print Cash flow c. How much cash flow ls lost due to the reduced depreciation from $40,000 to $20,0007 Cash tow lost
Answer #1

1)

Earnings before depreciation and taxes = 100,000

Depreciation = ( 40,000)

Earnings before taxes = 60,000

Taxes = 0.24 * 60,000 = (14,400)

Earnings after taxes = 45,600

Depreciation = 40,000

Cash flow = 85,600

2)

If dep was only 20,000:

Earnings before depreciation and taxes = 100,000

Depreciation = ( 20,000)

Earnings before taxes = 80,000

Taxes = 0.24 * 80,000 = (19,200)

Earnings after taxes = 60,800

Depreciation = 20,000

Cash flow = 80,800

c)

Cash flow lost = 85,600 - 80,800

Cash flow lost = 4,800